OpenStudy (anonymous):

What is the similarity ratio of the smaller to the larger similar pyramids? Enter your answer as a:b

OpenStudy (anonymous):



OpenStudy (anonymous):

ONE BIG ONE SMALL

Directrix (directrix):

Get the ratio of corresponding sides. 9/12 = 18/24 = ? when simplified

OpenStudy (anonymous):

ONE GOT MORE LENGHT TO IT

OpenStudy (anonymous):

I don't get it, do you cross multiply? @Directrix

Directrix (directrix):

What does 9/12 simplify to? Both 9 and 12 have a common factor of 3?

OpenStudy (anonymous):

3/4

Directrix (directrix):

ok 18/24 = ? Simplify. Both have common factors

OpenStudy (anonymous):

3/4

Directrix (directrix):

Yes. Similar solids have corresponding sides with the same ratio. That is the similarity ratio. You can write 3/4 as 3:4. Read it as: three to four.

OpenStudy (anonymous):

 alright, but what would you do with this one because it has three numbers?